OpenStudy (anonymous):

Does anyone knows a really good book where I can learn about logarithms? I want a lot of theory and a million exercises.

OpenStudy (anonymous):

logarithms arent fun

OpenStudy (anonymous):

I think the theory behind logarithms would lie in number theory

OpenStudy (anonymous):

also there is literally only like 4 things you need to any logarithm question

OpenStudy (anonymous):

you dont need "a million exercises"

OpenStudy (anonymous):

"logarithms arent fun" is that a book or is that what you think?

OpenStudy (anonymous):

yeh its a book :p

OpenStudy (anonymous):

Im trying to solve a lot of equations, inequalities, doing graphs, etc

OpenStudy (anonymous):

im out trying to promote my online tutoring....i tutor math and physics from low - hgh level....if u ever need it check me out http://mathphystut.tripod.com

OpenStudy (anonymous):

I know the theory behind logarithm in general, but i want something really challenging

OpenStudy (anonymous):

Solve \[2\ln(\sqrt{x}) - \ln(x) =1 \]

OpenStudy (anonymous):

an example

OpenStudy (anonymous):

but its all relatively standard

OpenStudy (anonymous):

You want challenging problems, or you want to clear your concept?

OpenStudy (anonymous):

want challenging problems

OpenStudy (anonymous):

yeh, well logarithms are that challenging to be honest

OpenStudy (anonymous):

arent*

OpenStudy (anonymous):

You can try one thing. You email me, and I will send you challenging problems

OpenStudy (anonymous):

As many as you want

OpenStudy (anonymous):

sure. can you give me your e-mail?

OpenStudy (anonymous):

About the problem above, elecengineer, I got 0=1. Is that right?

OpenStudy (anonymous):

Thanks "pi" I just sent you an email. Can you check it out

OpenStudy (anonymous):

It must take a little bit. Im from Ecuador and I suppose your in Europe lol

OpenStudy (anonymous):

I haven't yet received your email........... Still waiting

OpenStudy (anonymous):

Can you check it out again. Can take a little bit

OpenStudy (anonymous):

differentiate \[y= \ln [\frac{ (x-3)^4 \sqrt{x} }{x+1} ]\]

OpenStudy (anonymous):

thats prob the hardest question on logarithms every, but its still really easy

OpenStudy (anonymous):

Okay code is VILLAMAGUA CONZA LUIS MIGUEL VILLAMAGUA CONZA LUIS MIGUEL

OpenStudy (anonymous):

yean. I got 4LN(x-3)+1/2(LN(x))-LN(X+) cannot remember how to differenciate logs lol

OpenStudy (anonymous):

just joking

OpenStudy (anonymous):

yeah it is me PI, thanks for your help

OpenStudy (anonymous):

I have sent you a problem, but remember, you are not supposed to use a calculator

OpenStudy (anonymous):

lols, what was it

OpenStudy (anonymous):

By the way, are you a staff of UTPL ?

OpenStudy (anonymous):

what was the question!

OpenStudy (anonymous):

yeah. From UTPL

OpenStudy (anonymous):

You teach there?

OpenStudy (anonymous):

No teaching, but a student. Well finishing my major

OpenStudy (anonymous):

solve simultaneously \[5^{x+y} = \frac{1}{5}\ and \[5^{3x+2y} =1\]

OpenStudy (anonymous):

\[5^{x+y} = \frac{1}{5} , 5^{3x+2y} =1 \]

OpenStudy (anonymous):

I just replied you first question PI. Is that all right?

OpenStudy (anonymous):

there fairly standard lol all of logarithms are fairly standard. I have no idea why people find them hard , all it is is 4formulas to remember

OpenStudy (anonymous):

I dont know much about inequalities with logs.

OpenStudy (anonymous):

yeh , they come up a tiny bit in some financial maths topics

OpenStudy (anonymous):

inequalities and logs dont come up alot there is however one thing which you do need to look out for when solving inequalities with logs

OpenStudy (anonymous):

best seen by an example

OpenStudy (anonymous):

When I was in high school logarithms seems to be so hard. Now I tried them again and they look so easy. Im confused. I thought I would never be done with logs

OpenStudy (anonymous):

solve (1/3)^n > 0.5

OpenStudy (anonymous):

solve that, and see what happens

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I will put money on you being wrong :P

OpenStudy (anonymous):

it looks fairly simple, but there is a trick in it that would catch alot of people

OpenStudy (anonymous):

please dont go. Im a little bit slow. I like this :)

OpenStudy (anonymous):

well ... have you done the question yet inequalities with logs behave the exact same as equations

OpenStudy (amistre64):

the bigger n is.... the .......

OpenStudy (anonymous):

This is what i got

OpenStudy (anonymous):

long way to go about things also you would need to use change of basic at the end if you actually wanted to get a number for it

OpenStudy (anonymous):

What is the other wat? Im excited lol

OpenStudy (anonymous):

what is the other way? sorry

OpenStudy (anonymous):

well the way I was taught was to take logarithms of both sides ( it doesnt matter what base, as long as you use the same on both sides ) so I will take ln of both sides so \[\ln [( \frac{1}{3})^n] = \ln(0.5) \]

OpenStudy (anonymous):

now \[\log_{a} x^r = r \log_{a} x \]

OpenStudy (anonymous):

yeah. I know where you go

OpenStudy (anonymous):

your bright man

OpenStudy (anonymous):

so then you get n ln(1/3) >0.5

OpenStudy (anonymous):

*so then you get n ln(1/3) > ln(0.5)

OpenStudy (anonymous):

\[n > \frac{\ln(0.5)}{\ln(\frac{1}{3} ) }\]

OpenStudy (anonymous):

there the problem people make!

OpenStudy (anonymous):

mm, I divide both sides by ln(1/3) , but am I allowed to do that? well , ln(1/3) is actually a negative number , so I must flip the inequality sign

OpenStudy (anonymous):

ln(a) is negative for all a in the interval 0<a<1

OpenStudy (anonymous):

Yeah I would have fallen there too. Can you email me some of this type of exercices.